Kelly seels 80 shares of stock she owns for a total of $491. If the stock was in two different companies, one selling at $5.50 a share and the other at $6.25 a share, how much of each did she sell?
At $5.50
At $6.25

Answer by RAY100(1637)

given, total =80 shares =$491
let x = # shares @ $5.50
let y= # shares @ $6.25
491=5.50*x+6.25 y
but x +y=80,,,,therefore x=(80-y)
491 = (5.5)(80-y) + (6.25)(y)
491= 440-5.5y+6.25y
51=.75y
y=68
and therefore x=80-68=12
check
491=12(5.50)+68(6.25)=66+425=491,,,,ok